if mass percent of X in XO2 was 80 percent, how much will be X mass percent in XO?
1
Expert's answer
2012-12-24T12:11:28-0500
XO2 is 100 % and X is 80,so 2O is 20 %
& 2 atoms of Oxygen has molecular weight 16+16=32& (and it is 20 %)
Now you can find molecular weight of X using the proportion:

32----------------20%
X-----------------80%

X=128

So mass percent of X in XO is

w=128/(128+16)*100 %=88.89%